OpenStudy (christos):

I don't know how to solve this. I don't even know what is the name of the process I need to follow to get to what someone would call a "solution" (x+1)^1/3 + 1/3 (x+1) ^-2/3 When x CAN'T be -1 Medal will be given + 1 great fan testimonial

OpenStudy (anonymous):

Are you trying to solve:\[(x + 1)^{\frac{ 1 }{ 3 }} + (1/3)(x + 1)^{\frac{ -2 }{ 3 }} = 0\] ?

OpenStudy (anonymous):

Put both terms over one denominator:\[\frac{ (x + 1) + \frac{ 1 }{ 3 } }{ (x + 1)^{\frac{ 2 }{ 3 }} }\]

OpenStudy (anonymous):

Set the above equal to 0. Then all you have to do is see where (x + 1) + 1/3 = 0 x + 4/3 = 0 x = -4/3

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @Christos

OpenStudy (anonymous):

All good now, @Christos ?